OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (gorv):

add 5 on each side n tell what u got??

OpenStudy (anonymous):

if you add 5 then its 7sin2x-14sinx+2=0??

OpenStudy (gorv):

on both side

OpenStudy (gorv):

on left side also

OpenStudy (anonymous):

i meant to put a 7 instead of a 2

OpenStudy (anonymous):

7sin2x-14sinx+7=0?

OpenStudy (gorv):

now take out 7 from each term

OpenStudy (anonymous):

sin2x-2sinx+1 ?

OpenStudy (gorv):

yeah u see some thing now ??

OpenStudy (gorv):

(sinx)^2-2*1*sinx+1^2?? ring a bell ??

OpenStudy (anonymous):

no it doesnt im not good at this

OpenStudy (gorv):

(a-b)^2=a^2-2*a*b+b^2 now ....remember it ??

OpenStudy (anonymous):

oh yeah i know that i eaquation so do i just put it like that?

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

yeah we got in our what on right side ....we need to put it in like left side

OpenStudy (gorv):

what is similar to a??

OpenStudy (anonymous):

sin2x?

OpenStudy (gorv):

sinx

OpenStudy (gorv):

sin2x is a^2

OpenStudy (gorv):

a=sinx and b= 1

OpenStudy (gorv):

(sinx-1)^2=0

OpenStudy (gorv):

which means sinx-1=0

OpenStudy (gorv):

sinx =1

OpenStudy (anonymous):

ok so (sinx-1)^2=sinx^2-2*sinx*1+1^2

OpenStudy (gorv):

yeah

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

so then what do i do?

OpenStudy (gorv):

sinx=1

OpenStudy (gorv):

solution= (2n-1)*pi/2

OpenStudy (gorv):

n=1,2,3,4.................

OpenStudy (anonymous):

so the answer we would get from that is (2n-1)*pi/2

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

yeah

OpenStudy (anonymous):

ok Thnaks!

OpenStudy (gorv):

and n=1,2,3,4........and so on